Help pls 😊
Triangular numbers are: 1,3,6,10....where the nth such number is n(n-1)/2. Show that the sum of two consecutive triangular numbers is a square number.

Answer by Alan3354(69443)

Triangular numbers are: 1,3,6,10....where the nth such number is n(n-1)/2. Show that the sum of two consecutive triangular numbers is a square number.
------------
n(n-1)/2 + (n+1)*n/2 = (n^2-n + n^2+n)/2
= 2n^2/2
= n^2